# Classifying and Addressing the Diversity of Errors in Retrieval-Augmented Generation Systems

### v1

@@ Abstract
Retrieval-augmented generation (RAG) is a prevalent approach for building LLM-based question-answering systems that can take advantage of external knowledge databases. Due to the complexity of real-world RAG systems, there are many potential causes for erroneous outputs. Understanding the range of errors that can occur in practice is crucial for robust deployment. We present a new taxonomy of the error types that can occur in realistic RAG systems, examples of each, and practical advice for addressing them. Additionally, we curate a dataset of erroneous RAG responses annotated by error types. We then propose an auto-evaluation method aligned with our taxonomy that can be used in practice to track and address errors during development. Code and data are available at github.com/layer6ai-labs/rag-error-classification.
@@
Kin Kwan Leung Mouloud Belbahri Yi Sui Alex Labach Xueying Zhang Stephen Rose Jesse C. Cresswell Layer 6 AI, Toronto, Canada
@@ 3 RAG Implementation
Datasets: We use the DragonBall dataset, part of the RAGEval framework (Zhu et al., 2024), an evaluation suite designed to rigorously assess RAG systems across diverse domains and scenarios. The dataset provides a broad benchmark for evaluating RAG performance in complex settings, spanning domains such as finance, law, and medicine. We used both the English and Chinese portions of the dataset with 3108 and 3601 questions respectively. We also use the CLAPnq dataset for additional analysis (Rosenthal et al., 2025), which contains high-quality reference answers and passages for 4946 natural questions relating to Wikipedia documents.
@@ 4.3 Reranking
Improvements: Modern RAG pipelines increasingly include a reranking stage to reorder retrieved chunks based on semantic relevance Bleiweiss (2024). The two dominant approaches are cross-encoders, which jointly encode query-chunk pairs Déjean et al. (2024), and LLM-based rerankers, which use prompted models to assess candidates Sun et al. (2023). Rerankers may still miss key context (E7) or over-rank irrelevant chunks (E8) when domain relevance is poorly understood. Fine-tuning on in-domain data is the most effective way to improve performance Moreira et al. (2024); Aarsen (2025); Yu et al. (2024b), and even small amounts of labeled data can yield significant gains Aarsen (2025). More specifically, incorporating hard negatives, which are plausible but non-relevant passages, further helps the model reduce false positives (E8) Yu et al. (2024b).
@@ 5 RAGEC: RAG Error Classification
Data Curation. No existing dataset details the types of errors which can occur in RAG systems. Hence, to validate RAGEC, we manually annotated 406 erroneous responses generated by our reference system on the DragonBall dataset as to the responsible stage, and error types present. Annotators used all available context including ground truth answer, generated response, retrieved and reranked chunks, and the corpus. First, annotators selected the earliest stage in the pipeline which demonstrated erroneous behaviour, since errors propagate through RAG systems. Subsequently, one or more error types were selected within that stage. The repo at github.com/layer6ai-labs/rag-error-classification contains our annotated data.
Figure 2: Mode error type distribution per dataset.
@@
For error type classification, we query the LLM $K=10$ times for each datapoint and extract the most and second-most frequently predicted error types. Compared to human-annotated error types, RAGEC achieves an accuracy of 40.3%. Figure˜2 shows the distribution of the modal error type for DragonBall-EN and CLAPnq. Table˜2 shows more details for the DragonBall-EN dataset (CLAPnq is shown in Table˜5 of Section˜C.1). These distributions differ, showing that the nature of the dataset can affect the types of errors committed by a given RAG pipeline. This type of information can direct diagnosis and triaging of problems in RAG. For example, CLAPnq exhibited no chunking errors, as its source documents were manually curated by the dataset annotators into self-contained paragraphs designed to answer the corresponding queries. In contrast, DragonBall-EN almost never shows Parametric Overreliance (E11) for both English and Chinese (See Appendix C.2 for details). This can be attributed to the dataset’s domain-specific and context-dependent queries, which are unlikely to be covered by the LLM’s parametric knowledge. RAGEC also demonstrates that some error types dominate their stage, like E3 for chunking, while others like E5 primarily occur as a secondary error along with another type. Notably, Fabricated Content (E10)—often referred to as hallucination—is relatively rare in our analysis. This observation stands in contrast to the focus on hallucination in recent research and public discourse surrounding LLMs Ji et al. (2023). Our findings suggest that, in the context of RAG, other types of errors such as retrieval or chunking issues are more prevalent and may warrant greater attention. Finally, in Table˜3 we examine the consistency of classifications via mode frequency, the number of the $K$ runs that agreed with the mode error type. For the majority of the queries, RAGEC is quite consistent in determining error categories. Section˜C.3 discusses the co-occurrence of the error categories.
@@ Appendix A RAG Pipeline and Dataset Details
This section outlines the configurable components of our RAG pipeline and provides details on the specific configurations used to obtain the experimental results presented in the main paper. In Table˜4, we list important hyperparameters used on the DragonBall and CLAPnq datasets. Note that the DragonBall dataset is publicly available under a CC BY-NC-SA 4.0 license. The CLAPnq dataset is publicly available under an Apache 2.0 license.
@@ C.2 DragonBall-CN Error Type Distribution
To demonstrate that our method extends to languages other than English, we tested our method on the Chinese subset of the DragonBall dataset. Note that the documents and the queries are in Simplified Chinese. As in Table˜4, the parameters used for chunking, embedding, retrieval, reranking, and generation are all the same as DragonBall-EN, with the exception that chunking uses 512 characters instead of 128 tokens, which is more suitable for the structure of Chinese. We also express the generator prompt in Chinese, as listed below:

### v2

@@ Abstract
Retrieval-augmented generation (RAG) is a prevalent approach for building LLM-based question-answering systems that can take advantage of external knowledge databases. Due to the complexity of real-world RAG systems, there are many potential causes for erroneous outputs. Understanding the range of errors that can occur in practice is crucial for robust deployment. We present a new taxonomy of the error types that can occur in realistic RAG systems, examples of each, and practical advice for addressing them. Additionally, we curate a dataset of erroneous RAG responses annotated by error types. We then propose an auto-evaluation method aligned with our taxonomy that can be used in practice to track and address errors during development. Code and data are available at https://github.com/layer6ai-labs/rag-error-classification.
@@
Kin Kwan Leung Mouloud Belbahri Yi Sui Alex Labach Xueying Zhang Stephen Anthony Rose Jesse C. Cresswell Layer 6 AI, Toronto, Canada {kk, mouloud, amy, alex, chloe, stephen, jesse}@layer6.ai
@@ 3 RAG Implementation
Datasets: We use the DragonBall dataset, part of the RAGEval framework (Zhu et al., 2024), an evaluation suite designed to rigorously assess RAG systems across diverse domains and scenarios. The dataset provides a broad benchmark for evaluating RAG performance in complex settings, spanning domains such as finance, law, and medicine. The dataset also comes with query type annotations in seven categories, which allows us to perform thorough analysis on error classification. We used both the English and Chinese portions of the dataset with 3108 and 3601 questions respectively. We also use the CLAPnq dataset for additional analysis (Rosenthal et al., 2025), which contains high-quality reference answers and passages for 4946 natural questions relating to Wikipedia documents.
@@ 4.3 Reranking
Improvements: Modern RAG pipelines increasingly include a reranking stage to reorder retrieved chunks based on semantic relevance Bleiweiss (2024). The two dominant approaches are cross-encoders, which jointly encode query-chunk pairs déjean2024thorough, and LLM-based rerankers, which use prompted models to assess candidates Sun et al. (2023). Rerankers may still miss key context (E7) or over-rank irrelevant chunks (E8) when domain relevance is poorly understood. Fine-tuning on in-domain data is the most effective way to improve performance Moreira et al. (2024); Aarsen (2025); Yu et al. (2024b), and even small amounts of labeled data can yield significant gains Aarsen (2025). More specifically, incorporating hard negatives, which are plausible but non-relevant passages, further helps the model reduce false positives (E8) Yu et al. (2024b).
@@ 5 RAGEC: RAG Error Classification
Data Curation. No existing dataset details the types of errors which can occur in RAG systems. Hence, to validate RAGEC, we manually annotated 406 erroneous responses generated by our reference system on the DragonBall dataset as to the responsible stage, and error types present. Annotators used all available context including ground truth answer, generated response, retrieved and reranked chunks, and the corpus. First, annotators selected the earliest stage in the pipeline which demonstrated erroneous behaviour, since errors propagate through RAG systems. Subsequently, one or more error types were selected within that stage. The repo at https://github.com/layer6ai-labs/rag-error-classification contains our annotated data.
Figure 2: Modal error type distribution per dataset.
@@
For error type classification, we query the LLM $K=10$ times for each datapoint and extract the most and second-most frequently predicted error types. Compared to human-annotated error types, RAGEC achieves an accuracy of 40.3%. Figure˜2 shows the distribution of the modal error type for DragonBall-EN and CLAPnq. Table˜2 shows more details for the DragonBall-EN dataset (CLAPnq is shown in Table˜6 of Section˜C.1). These distributions differ, showing that the nature of the dataset can affect the types of errors committed by a given RAG pipeline. This type of information can direct diagnosis and triaging of problems in RAG. For example, CLAPnq exhibited no chunking errors, as its source documents were manually curated by the dataset annotators into self-contained paragraphs designed to answer the corresponding queries. In contrast, DragonBall-EN almost never shows Parametric Overreliance (E11) for both English and Chinese (See Appendix C.2 for details). This can be attributed to the dataset’s domain-specific and context-dependent queries, which are unlikely to be covered by the LLM’s parametric knowledge. RAGEC also demonstrates that some error types dominate their stage, like E3 for chunking, while others like E5 primarily occur as a secondary error along with another type. Notably, Fabricated Content (E10)—often referred to as hallucination—is relatively rare in our analysis. This observation stands in contrast to the focus on hallucination in recent research and public discourse surrounding LLMs Ji et al. (2023). Our findings suggest that, in the context of RAG, other types of errors such as retrieval or chunking issues are more prevalent and may warrant greater attention.
@@
Next we demonstrate how insights can be gathered into a RAG pipeline’s behaviour through fine-grained analysis with RAGEC. Table˜3 shows the distribution of the error types across knowledge domains and query types defined in DragonBall-EN.
Error Types across Knowledge Domains. We see that finance queries lead to the most diversity in error types, but also skew more towards generation errors. Misinterpretation (E13) is the single biggest challenge. For law queries three of the RAG stages contribute roughly the same amount of errors, and overall law is the most difficult domain. Numerical errors (E16) are more common for law than any other domain which may reveal an underlying deficiency in the generation model around calculations that are typically done in law, as opposed to finance. Medical queries are by far dominated by chunking errors, with context mismatch (E3) having an outsized presence. A practitioner seeing this could simply increase chunk size, return entire documents corresponding to the top chunks, or invest in more complex solutions like expanding chunked text to include explicit definitions.
Error Types across Query Types. For factual questions three error types dominate (E3, E4, E13), whereas unanswerable questions, where the correct response is to abstain, show the highest diversity in error types. As expected, the abstention failure (E9) error type is used nearly exclusively for unanswerable questions. Summarization queries, the most difficult type, are dominated by chunking errors (E2, E3). Multiple semantically related passages are typically needed to produce a coherent and complete summary, and these connections can be lost because of chunking. Numerical comparison queries show fewer generation errors than other types, and numerical errors (E16) are not particularly troublesome. Errors are instead coming from context mismatch (E3) and missed retrieval (E4). Finally, as expected, multi-hop reasoning is one of the more challenging types and displays a wide range of error types for this RAG implementation. This range accurately reflects the more complicated nature of multi-hop reasoning queries where many things can go wrong, and demonstrates that the proposed error categories are expressive enough to capture the diversity of errors in complex query types.
Table 3: Modal error type distributions for DragonBall-EN across knowledge domains and query types. Query type labels are provided by the dataset.
Error Classification Knowledge Domain Query Type
Finance Law Medicine
Factual
Question
Unanswerable
Question
Numerical
Comparison
Information
Integration
Temporal
Sequence
Multi-hop
Reasoning
Summari-
zation
E1 Overchunking 0 0 0 0 0 0 0 0 0 0
E2 Underchunking 2 19 11 1 3 1 0 1 7 19
E3 Context Mismatch 19 71 117 19 5 22 43 12 24 82
Percentage Chunking 5.9% 32.8% 62.4% 36.4% 16.3% 29.1% 18% 14.6% 27% 49%
E4 Missed Retrieval 94 76 59 22 7 29 87 35 20 29
E5 Low Relevance 25 6 2 1 0 4 20 1 3 4
E6 Semantic Drift 0 0 0 0 0 0 0 0 0 0
Percentage Retrieval 33.7% 29.9% 29.8% 41.8% 14.3% 41.8% 44.8% 40.4% 20% 16%
E7 Low Recall 28 23 2 0 2 3 25 10 8 5
E8 Low Precision 4 2 0 1 1 2 0 1 0 1
Percentage Reranking 9.1% 9.1% 1% 1.8% 6.1% 6.3% 10.5% 12.4% 7% 2.9%
E9 Abstention Failure 5 1 0 0 5 0 1 0 0 0
E10 Fabricated Content 30 4 0 0 4 2 5 0 8 15
E11 Parametric Overreliance 4 0 0 0 0 0 0 0 1 3
E12 Incomplete Answer 8 8 1 0 0 0 7 0 8 2
E13 Misinterpretation 116 50 12 10 16 11 46 29 26 40
E14 Contextual Misalignment 11 2 1 0 4 2 3 0 0 5
E15 Chronological Inconsistency 2 1 0 1 0 0 1 0 0 1
E16 Numerical Error 5 11 0 0 2 3 1 0 10 0
Percentage Generation 51.3% 28.1% 6.8% 20% 63.3% 22.8% 26.8% 32.6% 46.1% 32%
Total Errors 353 274 205 55 49 79 239 89 115 206
Total Queries 1574 782 752 535 233 297 737 241 532 533
Error Rates 22.4% 35% 27.3% 10.3% 21% 26.6% 32.4% 36.9% 21.6% 38.6%
@@
Classification Consistency. Finally, in Table˜4 we examine the consistency of classifications via mode frequency, the number of the $K$ runs that agreed with the modal error type. For the majority of the queries, RAGEC is quite consistent in determining error categories. Section˜C.3 discusses the co-occurrence of the error categories.
@@ Appendix A RAG Pipeline and Dataset Details
This section outlines the configurable components of our RAG pipeline and provides details on the specific configurations used to obtain the experimental results presented in the main paper. In Table˜5, we list important hyperparameters used on the DragonBall and CLAPnq datasets. Note that the DragonBall dataset is publicly available under a CC BY-NC-SA 4.0 license. The CLAPnq dataset is publicly available under an Apache 2.0 license.
@@ C.2 DragonBall-CN Error Type Distribution
To demonstrate that our method extends to languages other than English, we tested our method on the Chinese subset of the DragonBall dataset. Note that the documents and the queries are in Simplified Chinese. As in Table˜5, the parameters used for chunking, embedding, retrieval, reranking, and generation are all the same as DragonBall-EN, with the exception that chunking uses 512 characters instead of 128 tokens, which is more suitable for the structure of Chinese. We also express the generator prompt in Chinese, as listed below:
@@ C.4 Detailed Error Classification Results
Appendix D AI assistant usage
In this research, we used ChatGPT to assist paper editing, literature search, and Github Copilot for coding assistance.